# Coupled Cluster Downfolding Theory in Simulations of Chemical Systems on Quantum Hardware

### v1

@@ I Introduction
The accurate characterization of the complex correlated behavior of electrons in molecules emerges as one of the most pressing challenges in the many-body theory of interacting systems. Several classes of methodologies, based on various representations of quantum mechanics including wave function Coester (1958); Čížek (1966); Raghavachari et al. (1989); Bartlett and Musiał (2007), electron density Lee et al. (1988); Perdew et al. (1996); Jones (2015), density matrix Knizia and Chan (2012); Wouters et al. (2016), and Green’s function Hedin (1965); Aryasetiawan and Gunnarsson (1998); Van Setten et al. (2015); Rehr et al. (2010) based methods have been introduced, tested, and validated to address this problems. Significant progress has been achieved in density functional theory, density matrix renormalization group (DMRG) Menczer et al. (2024), coupled cluster (CC) formalisms Haugland et al. (2020); Yuwono et al. (2024); Windom et al. (2024); Gururangan et al. (2025), as well as variational formulations Tubman et al. (2016); Schriber and Evangelista (2016); Abraham and Mayhall (2020, 2022) and self-energy-based approaches van Schilfgaarde et al. (2006); Deslippe et al. (2012); Reeves et al. (2025). Applications of these methods go well beyond quantum chemistry - analogous formulations are currently used in the context of materials sciences and nuclear physics (see, for example, Ref.26; 27; 28; 29; 30).
The high-accuracy results generated by hierarchical classes of approximations, i.e., sequences of approximations that can reach the exact diagonalization limit, are usually associated with significant numerical overhead as a function of the system size as well as memory requirements needed to include parameter definitions, for example, wave function expansions. To alleviate these roadblocks of conventional computing, several strategies to address these bottlenecks have been developed over the last few decades. These include the development of a scalable library for parallel implementations of electronic structure methods and a broad class of reduced-scaling approaches based on the local character of correlation effects. In recent years, several formulations have been introduced to effectively utilize the sparsity/locality of the general quantum systems and various types of downfolding or embedding procedures. However, pushing the system-size limit and maintaining the required level of accuracy using these two strategies may come up against natural limits, for example, due to the size-extensive character of the errors made in the so-called local formulations or incomplete characterization of the interactions between embedded region/regions and the surrounding environment.
@@
Quantum computing provides an alternative computational model to bypass the typical challenges of conventional computing associated with speed, memory, and energy limitations needed to achieve the so-called chemical accuracy. Algorithms such as Quantum Phase Estimation (QPE) have the potential to overcome the exponential computational barriers preventing classical computing from reaching the exact diagonalization limit in simulations for realistic systems. However, for this to happen, the quantum hardware must mature to the point where deep quantum circuits and quantum error correction can be effectively implemented at scale. Given the existing limitations of the quantum hardware, the hardware realization of QPE-type algorithms is scarce and still limited to the small-size, toy systems Aspuru-Guzik et al. (2005); Wecker et al. (2014); Wiebe and Granade (2016); Santagati et al. (2018); McArdle et al. (2020); Klymko et al. (2022). For this reason, simpler and less efficient methods, in the context of the correlation effects included, are used. An excellent example is provided by the class of hybrid Variational Quantum Eigensolvers (VQE). Although hardware VQE executions for small chemical systems have been reported, their applications are also limited by circuit complexity and the need for explicit inclusion of a large number of wave function parameters, which makes VQE approaches not scalable. Although hybrid VQE applications can be performed for larger systems than possible with the QPE approach, their applications are limited to small-size active spaces, which in the best-case scenario can only encapsulate part of the static correlation effects. However, without an efficient way of incorporating dynamical correlation effects, usually associated with virtual orbitals falling outside of the active spaces, the accuracy of both QPE and VQE formalisms falls short of recovering the total correlation energies comprising both static and dynamical components needed to obtain an adequate level of accuracy in simulations for chemical processes.
@@
A practical approach for tackling these challenges while simultaneously providing a strategy for utilizing the resources of current quantum computers is linked to recent advancements in various dimensionality-reduction techniques of quantum problems, hereafter referred to as downfolding formalism Bauman et al. (2019); Kowalski and Bauman (2020); Huang et al. (2023). These formulations lay the foundation for efficient hybrid quantum/classical execution protocols. For example, the recently introduced coupled cluster downfolding formalisms allow the construction of effective Hamiltonians that encapsulate dynamical correlation effects within the active spaces whose dimensions can be tuned to the available quantum resources. The CC downfolding introduced in Refs.40; 37; 38; 41; 39; 42 or other embedding techniques based on the CC downfolding techniques Peng and Kowalski (2024); Shee et al. (2024) utilize the simplicity of single-reference unitary CC methods and provide a platform for hierarchical inclusion of collective many-body effects. The CC downfolding method, in conjunction with quantum solvers based on the exponential form of the wave function ansatz (i.e. unitary CC formulations Hoffmann and Simons (1988); Bartlett et al. (1989); Lee et al. (2018)), in contrast to quantum algorithms targeting some form of truncated configuration interaction formalism, provides size-consistency or size-extensivity of calculated energies (i.e., additive separability of energy in the noninteracting subsystem limit), a property that has been identified by John Pople in his Nobel Lecture Pople (1999) as a matter of great importance. We argue that satisfying this requirement should be addressed at the early stages of applications of quantum computing in chemistry. Furthermore, its multi-active space variant, referred to as the Quantum Flow approach (QFlow) Kowalski and Bauman (2023, 2025), facilitates the exploration of extensive subspaces within Hilbert space through the use of reduced quantum resources and constant-depth quantum circuits, giving rise to efficient utilization of distributed quantum resources.
@@ II Results
  • Free-base porphyrin: The C20H14N4 system, in analogy to the benzene case, is characterized by the presence of strong dynamical correlation effects amounting to nearly 3 Hartree of correlation energy. As for the benzene molecule, the correlation energy depends almost linearly on the number of correlated virtual orbitals (see Fig. 2(b)).
In our simulations, we employed Cartesian geometries for benzene and FBP used in Refs. Eriksen et al., 2020 and Kowalski et al., 2010, respectively. For N2 the equilibrium geometry Re is assumed to be 2.068 $a_{0}$. In N2 simulations, we correlated all molecular orbitals, while in simulations for the benzene and FBP systems, we kept all core orbitals frozen. Table 1 presents ground-state energies obtained using several quantum chemical methods implemented on classical computers, including restricted Hartree–Fock (RHF), CCSD Purvis and Bartlett (1982), CCSD(T) Raghavachari et al. (1989), and CCSDTQ Kucharski and Bartlett (1991); Oliphant and Adamowicz (1991a); Piecuch and Adamowicz (1994), to evaluate the accuracy of the results from quantum hardware. For the benzene molecule in the cc-pVDZ basis set, a comparison of the CCSDTQ energies with other formulations indicates that the CCSD(T) approach accurately captures the non-dynamical correlation energy, which is the dominant contribution to the total correlation energy.
@@
The effective Hamiltonians are constructed in active spaces defined by the RHF orbitals. For practical reasons, primarily due to the limitations of quantum hardware and the associated cost of accessing these systems, we opted to use relatively small active spaces, defined by six active electrons distributed over six active orbitals for all benchmark systems, i.e., ($6e,6o$) active spaces. The hybrid execution workflow comprises classical computations to generate effective Hamiltonians, followed by energy evaluation and optimization on a quantum simulator or hardware.
The energies evaluated on the noiseless simulators are presented in Table 3, where they are compared with full configuration interaction (FCI) results obtained via exact diagonalization of downfolded Hamiltonians in ($6e,6o$) active spaces. For larger systems dominated by dynamical correlation effects (i.e., benzene and FBP), we observe excellent agreement between the energies computed using all noiseless simulators and the corresponding FCI values. A similar level of agreement is seen for the N2 system at its equilibrium geometry (RN-N = 1.0Re). For larger bond stretches of the N–N distance, where static correlation effects become increasingly significant, we find that among all noiseless simulators, the ADAPT-GCIM(2,2) formulation of Ref. 52 and the UCCGSD approach provide the closest agreement with the FCI results.
Given the limited quantum resources, we focused exclusively on simulations of the benzene and FBP systems. The ground-state energies obtained from hardware experiments for benzene (using the cc-pVTZ basis set) and FBP (using the cc-pVDZ basis set) are presented in Table 3. In all experiments, we employed downfolded Hamiltonians based on the double unitary coupled cluster ansatz (DUCC) consistent with the DUCC(3)-A(7) approximation described in Ref.61, where the $\sigma_{\rm ext}$-dependent components include single, double, and lowest-order triple commutators (see Appendix A for further details). For both systems, the energies measured on the Quantinuum H1-1 quantum computer show a good agreement with those obtained using the H1-1 Emulator for non-noise-amplified circuits. Specifically, the energy discrepancy for benzene is $3.4$ milliHartree, while for FBP it is $6.7$ milliHartree.
@@
C6H6 cc-pVDZ 114 -230.7218 -231.5453 -231.5813 -231.5842111CCSDTQ energy taken from Ref.55.
@@ III.1 CC downfolding and effective Hamiltonians
Coupled-cluster (CC) downfolding techniques have emerged as powerful tools for reducing the dimensionality of quantum many-body problems. Central to these approaches is the exponential CC ansatz, which facilitates the construction of effective (downfolded) Hamiltonians within a reduced-dimensional active space. These effective Hamiltonians encode the influence of out-of-active-space correlation effects (i.e., dynamical correlation) into their many-body structure. This formulation ensures that the lowest eigenvalue of the effective Hamiltonian corresponds to the exact or approximate CC energy.
Two primary variants of CC downfolding exist, based on the standard single-reference CC and the more general unitary CC ansatz. The latter leads to Hermitian CC downfolding, which is particularly relevant for quantum computing applications, although any Hermitian solver can be used for diagonalization. In contrast, non-Hermitian CC downfolding provides a mathematically elegant framework for connecting different active-space problems, an idea formalized in the Equivalence Theorem (see Ref. 41). This connection underpins the concept of quantum flow Kowalski and Bauman (2023), which is also explored within the Hermitian formalism.
@@
The Hermitian formulations utilize the DUCC ansatzBauman et al. (2019); Kowalski and Bauman (2020), where the exact ground-state wave function is represented as the product of two exponential ansatzes defined by anti-Hermitian cluster operators $\sigma_{\rm ext}$ (similar decomposition of the standard cluster operator is employed in active-space CC formulations see Refs.82; 83; 84 ) and $\sigma_{\rm int}$ that are expressed in terms of tensor parameters containing all spin-orbitals indices active and at least one inactive and no inactive spin-orbital indices, respectively. Using DUCC ansatz allows one can build (once certain conditions are met) an effective Hamiltonian $H^{\rm eff}$ in the active space
@@
which can reproduce the lowest eigenvalue being exact (or approximate) energy $E$ ($E^{\prime}$) of the system once exact (approximate) form of $\sigma_{\rm ext}(\mathfrak{h})$ is known. The $P+Q_{\rm int}$ is a projection operator onto active space ($P=|\Phi\rangle\langle\Phi|$ is a projection operator onto the reference function $|\Phi\rangle$ and $Q_{\rm int}$ represents a projection operator onto all excited configurations with respect to $|\Phi\rangle$ in active space). Using commutator expansion for Eq. (1) one can show that the $H^{\rm eff}$ operator is expressed in terms of connected diagrams only.
In general, due to the non-commutativity of the many-body components that define $\sigma_{\rm ext}$, analyzing the structure of the Hermitian effective Hamiltonian is more challenging than in the non-Hermitian case. Practical implementations require several approximations, including the treatment of the similarity transformation (see Eq. (1)), the representation of the $\sigma_{\rm ext}$ operator, and the truncation of many-body effects in the effective Hamiltonian. These challenges are typically addressed through: (i) a finite commutator expansion based on the Campbell–Hausdorff theorem; (ii) an approximation of the $\sigma_{\rm ext}$ operator using the unitary coupled cluster formalism, which leverages cluster amplitudes obtained from CCSD simulations; and (iii) the restriction of the effective Hamiltonian to one- and two-body terms.
@@ III.2 CC downfolding codes
The development of coupled-cluster (CC) downfolding computational infrastructure has been the focus of intensive efforts to establish an HPC framework capable of leveraging massively parallel GPU-based architectures for tackling real-world chemistry problems. Initial pilot implementations—hand-coded and serial—were instrumental in defining the hierarchical structure of the downfolded Hamiltonians based on double unitary coupled-cluster (DUCC) approximations, including classes of DUCC(2) and DUCC(3) approximate models Bauman and Kowalski (2022) (see also Appendix A). These early implementations assumed that all occupied orbitals were active Metcalf et al. (2020); Nicholas P et al. (2021).
@@ III.3 Database of downfolded Hamiltonians
The effective Hamiltonians of the molecular systems in this study are hosted on a public library repository (https://github.com/npbauman/DUCC-Hamiltonian-Library.git). The downfolded Hamiltonian library contains other molecular systems with varying active space sizes, different DUCC approximations, and different parent basis sets used to source the external CC amplitudes. Each system contains the ExaChem input and outputs and the fermionic effective Hamiltonians in various file formats, including raw text, FCIDUMP, and YAML based on the Broombridge schemaLow et al. (2019). Additional information about the effective Hamiltonians can also be found in the library, such as orbital energies and corresponding RHF and FCI energies for easy comparison and validation.

### v3

@@ I Introduction
The accurate characterization of the complex correlated behavior of electrons in molecules emerges as one of the most pressing challenges in the many-body theory of interacting systems. Several classes of methodologies, based on various representations of quantum mechanics including wave function Coester (1958); Čížek (1966); Raghavachari et al. (1989); Bartlett and Musiał (2007), electron density Lee et al. (1988); Perdew et al. (1996); Jones (2015), density matrix Knizia and Chan (2012); Wouters et al. (2016), and Green’s function Hedin (1965); Aryasetiawan and Gunnarsson (1998); Van Setten et al. (2015); Rehr et al. (2010) based methods have been introduced, tested, and validated to address this problems. Significant progress has been achieved in density functional theory Teale et al. (2022), density matrix renormalization group (DMRG) Menczer et al. (2024), coupled cluster (CC) formalisms Haugland et al. (2020); Yuwono et al. (2024); Windom et al. (2024); Gururangan et al. (2025), as well as variational formulations Tubman et al. (2016); Schriber and Evangelista (2016); Abraham and Mayhall (2020, 2022) and self-energy-based approaches van Schilfgaarde et al. (2006); Deslippe et al. (2012); Reeves et al. (2025). Applications of these methods go well beyond quantum chemistry - analogous formulations are currently used in the context of materials sciences and nuclear physics (see, for example, Refs.27; 28; 29; 30; 31).
@@
High-accuracy results generated by hierarchical classes of approximations -i.e., sequences of approximations that can reach the exact diagonalization limit - are usually associated with significant computational overhead and memory requirements as the system size increases, due to the large number of parameters needed to define accurate wave function expansions. Quantum computing provides an alternative computational model to bypass the typical challenges of conventional computing associated with speed, memory, and energy limitations needed to achieve the so-called chemical accuracy. Algorithms such as Quantum Phase Estimation (QPE) Nielsen and Chuang (2011); Kitaev (1995, 1997); Abrams and Lloyd (1999); Childs (2010); Reiher et al. (2017) have theoretically the potential to overcome the exponential computational barriers preventing classical computing from reaching the exact diagonalization limit in simulations for realistic systems. However, for this to happen, the quantum hardware must mature to the point where deep quantum circuits and quantum error correction can be effectively implemented at scale. Given the existing limitations of the quantum hardware, the hardware realization of QPE-type algorithms is scarce and still limited to the small-size, toy systems Aspuru-Guzik et al. (2005); Wecker et al. (2014); Wiebe and Granade (2016); Santagati et al. (2018); McArdle et al. (2020); Klymko et al. (2022). For this reason, simpler algorithms are currently being used. An excellent example is provided by the class of hybrid Variational Quantum Eigensolvers (VQE)Peruzzo et al. (2014); McClean et al. (2016); Romero et al. (2018); Shen et al. (2017); Kandala et al. (2017, 2019); Colless et al. (2018); Huggins et al. (2020); Cao et al. (2019); Weaving et al. (2025). Although hardware VQE executions for small chemical systems have been reported, their applications are also limited by circuit complexity and the need for explicit inclusion of a large number of wave function parameters, which makes VQE approaches not scalable. Although hybrid VQE applications can be performed for larger systems than possible with the QPE approach, their applications are limited to small-size active spaces, which in the best-case scenario can only encapsulate part of the static correlation effects. However, without an efficient way of incorporating dynamical correlation effects, usually associated with virtual orbitals falling outside of the active spaces, the accuracy of both QPE and VQE formalisms falls short of recovering the total correlation energies comprising both static and dynamical components needed to obtain an adequate level of accuracy in simulations for chemical processes.
@@
A practical approach for tackling these challenges while simultaneously providing a strategy for utilizing the resources of current quantum computers is linked to recent advancements in various dimensionality-reduction techniques of quantum problems, hereafter referred to as downfolding formalism Bauman et al. (2019); Kowalski and Bauman (2020); Huang et al. (2023). These formulations lay the foundation for efficient hybrid quantum/classical execution protocols. For example, the recently introduced coupled cluster downfolding formalisms allow the construction of effective Hamiltonians that encapsulate dynamical correlation effects within the active spaces whose dimensions can be tuned to the available quantum resources. The CC downfolding introduced in Refs.57; 54; 55; 58; 56; 59 or other embedding techniques based on the CC downfolding techniques Peng and Kowalski (2024); Shee et al. (2024) utilize the simplicity of single-reference unitary CC methods and provide a platform for hierarchical inclusion of collective many-body effects. The CC downfolding method, in conjunction with quantum solvers based on the exponential form of the wave function ansatz (i.e. unitary CC formulations Hoffmann and Simons (1988); Bartlett et al. (1989); Lee et al. (2018)), in contrast to quantum algorithms targeting some form of truncated configuration interaction formalism, provides size-consistency or size-extensivity of calculated energies (i.e., additive separability of energy in the noninteracting subsystem limit), a property that has been identified by John Pople in his Nobel Lecture Pople (1999) as a matter of great importance. We argue that satisfying this requirement should be addressed at the early stages of applications of quantum computing in chemistry. Furthermore, its multi-active space variant, referred to as the Quantum Flow approach (QFlow) Kowalski and Bauman (2023, 2025), facilitates the exploration of extensive subspaces within Hilbert space through the use of reduced quantum resources and constant-depth quantum circuits, giving rise to efficient utilization of distributed quantum resources. Some QFlow formulations require no prior knowledge about the structure of the sought-after state.
@@ II Results
  • Free-base porphyrin: The C20H14N4 system, in analogy to the benzene case, is characterized by the presence of strong dynamical correlation effects amounting to nearly 3.5 Hartree of correlation energy. As for the benzene molecule, the correlation energy depends almost linearly on the number of correlated virtual orbitals (see Fig. 2(b)).
In our simulations, we employed Cartesian geometries for benzene and FBP used in Refs. Eriksen et al., 2020 and Kowalski et al., 2010, respectively. For N2 the equilibrium geometry Re is assumed to be 2.068 $a_{0}$. In N2 simulations, we correlated all molecular orbitals, while in simulations for the benzene and FBP systems, we kept all core orbitals frozen. Table 1 presents ground-state energies obtained using several quantum chemical methods implemented on classical computers, including restricted Hartree–Fock (RHF), CCSD Purvis and Bartlett (1982), CCSD(T) Raghavachari et al. (1989), and CCSDTQ Kucharski and Bartlett (1991); Oliphant and Adamowicz (1991a); Piecuch and Adamowicz (1994) approaches, to evaluate the accuracy of the results from quantum hardware. For the benzene molecule in the cc-pVDZ basis set, a comparison of the CCSDTQ energies with other formulations indicates that the CCSD(T) approach accurately captures the non-dynamical correlation energy, which is the dominant contribution to the total correlation energy.
@@
The effective Hamiltonians are constructed in active spaces defined by the RHF orbitals. For practical reasons, primarily due to the limitations of quantum hardware and the associated cost of accessing these systems, we opted to use relatively small active spaces, defined by six active electrons distributed over six active orbitals for all benchmark systems, i.e., ($6e,6o$) active spaces. The hybrid execution workflow comprises classical computations to generate effective Hamiltonians, followed by energy evaluation and optimization on quantum simulators or quantum hardware.
The energies evaluated on the noiseless simulators are presented in Table 3, where they are compared with full configuration interaction (FCI) results obtained via exact diagonalization of downfolded Hamiltonians in ($6e,6o$) active spaces. For larger systems dominated by dynamical correlation effects (i.e., benzene and FBP), we observe excellent agreement between the energies computed using all noiseless simulators and the corresponding FCI values. A similar level of agreement is seen for the N2 system at its equilibrium geometry (RN-N = 1.0Re). For larger bond stretches of the N–N distance, where static correlation effects become increasingly significant, we find that among all noiseless simulators, the ADAPT-GCIM(2,2) formulation of Ref. 69 and the UCCGSD approach provide the closest agreement with the FCI results.
Given the limited quantum resources, we focused exclusively on hardware simulations of the benzene and FBP systems. The ground-state energies obtained from hardware experiments for benzene (using the cc-pVTZ basis set) and FBP (using the cc-pVDZ basis set) are presented in Table 3. In all experiments, we employed downfolded Hamiltonians based on the double unitary coupled cluster ansatz (DUCC) consistent with the DUCC(3)-A(7) approximation described in Ref.78 (see Appendix A for further details). For both systems, the energies measured on the Quantinuum H1-1 quantum computer show a good agreement with those obtained using the H1-1 Emulator for non-noise-amplified circuits. Specifically, the energy discrepancy for benzene is $3.4$ milliHartree, while for FBP it is $6.7$ milliHartree.
@@
C6H6 cc-pVDZ 114 -230.7218 -231.5453 -231.5813 -231.5842111CCSDTQ energy taken from Ref.72.
@@ III.1 CC downfolding and effective Hamiltonians
Coupled-cluster downfolding techniques have emerged as powerful tools for reducing the dimensionality of quantum many-body problems. Central to these approaches is the exponential CC ansatz, which facilitates the construction of effective (downfolded) Hamiltonians within a reduced-dimensional active space. These effective Hamiltonians encode the influence of out-of-active-space correlation effects (i.e., dynamical correlation) into their many-body structure. This formulation ensures that the lowest eigenvalue of the effective Hamiltonian corresponds to the exact or approximate CC energy.
Two primary variants of CC downfolding exist, based on the standard single-reference CC Kowalski (2018) and the more general unitary CC ansatz Bauman et al. (2019). The latter leads to Hermitian CC downfolding, which is particularly relevant for quantum computing applications. In contrast, non-Hermitian CC downfolding provides a mathematically elegant framework for connecting different active-space problems, an idea formalized in the Equivalence Theorem (see Ref. 58). This connection underpins the concept of Quantum Flow Kowalski and Bauman (2023), which is also explored within the Hermitian formalism.
@@
The Hermitian formulations utilize the DUCC ansatzBauman et al. (2019); Kowalski and Bauman (2020), where the exact ground-state wave function is represented as the product of two exponential ansatzes defined by anti-Hermitian cluster operators $\sigma_{\rm ext}$ (similar decomposition of the standard cluster operator is employed in active-space CC formulations see Refs.99; 100; 101 ) and $\sigma_{\rm int}$ that are expressed in terms of parameter containing all spin-orbital indices active ($\sigma_{\rm int}$) and at least one spin-orbital index inactive ($\sigma_{\rm ext}$). Using the DUCC ansatz, one can construct (once certain conditions are met) an effective Hamiltonian $H^{\rm eff}$ in the active space
@@
which can reproduce the lowest eigenvalue being exact (or approximate) energy $E$ ($E^{\prime}$) of the system once exact (approximate) form of $\sigma_{\rm ext}$ is known. The $P+Q_{\rm int}$ is a projection operator onto active space ($P=|\Phi\rangle\langle\Phi|$ is a projection operator onto the reference function $|\Phi\rangle$ and $Q_{\rm int}$ represents a projection operator onto all excited configurations with respect to $|\Phi\rangle$ in active space). Using commutator expansion for Eq. (1) one can show that the $H^{\rm eff}$ operator is expressed in terms of connected diagrams only.
In general, due to the non-commutativity of the many-body components that define $\sigma_{\rm ext}$, analyzing the structure of the Hermitian effective Hamiltonian is more challenging than in the non-Hermitian case. Practical implementations require several approximations, including the treatment of the similarity transformation (see Eq. (1)), the representation of the $\sigma_{\rm ext}$ operator, and the truncation of many-body effects in the effective Hamiltonian. These challenges are typically addressed through: (i) a finite commutator expansion based on the Campbell–Hausdorff theorem; (ii) an approximation of the $\sigma_{\rm ext}$ operator using the unitary coupled cluster formalism, which leverages cluster amplitudes obtained from the standard CCSD simulations; and (iii) the restriction of the effective Hamiltonian to one- and two-body terms.
@@ III.2 CC downfolding codes
The development of coupled-cluster downfolding computational infrastructure has been the focus of intensive efforts to establish an HPC framework capable of leveraging massively parallel GPU-based architectures for tackling real-world chemistry problems. Initial pilot implementations—hand-coded and serial—were instrumental in defining the hierarchical structure of the downfolded Hamiltonians based on double unitary coupled-cluster (DUCC) approximations, including classes of DUCC(2) and DUCC(3) approximate models Bauman and Kowalski (2022) (see also Appendix A). These early implementations assumed that all occupied orbitals were active Metcalf et al. (2020); Nicholas P et al. (2021).
@@ III.3 Database of downfolded Hamiltonians
The effective Hamiltonians of the molecular systems in this study are hosted on a public library repository (https://github.com/npbauman/DUCC-Hamiltonian-Library). The downfolded Hamiltonian library contains other molecular systems with varying active space sizes, different DUCC approximations, and different parent basis sets used to source the external CC amplitudes. Each system contains the ExaChem input and outputs and the fermionic effective Hamiltonians in various file formats, including raw text, FCIDUMP, and YAML based on the Broombridge schemaLow et al. (2019). Additional information about the effective Hamiltonians can also be found in the library, such as orbital energies and corresponding RHF and FCI energies for easy comparison and validation.